# Evaluating Anomaly Detectors for Simulated Highly Imbalanced Industrial Classification Problems

### v1

@@ Abstract
This work may be submitted for possible publication. Copyright may be transferred without notice.
@@ I-A4 Summary
Category Source Name Year
Number of
Examples
Anomaly
Rate
@@ II-B Available Datasets
Class
Distribution
Type
Parameters
Healthy
Three-mode
Gaussian
Mixture
$\sigma^{2}_{A}>0.01$, $\mu>0$,
$\mu_{A}=\sqrt{\mu^{2}/d}$,
$\mu_{1}=\vec{0}$,
$\mu_{2}=\mu_{A}\vec{1}$,
$\mu_{3}=-\mu_{A}\vec{1}$
Faulty
Gaussian
Mixture
Hypersphere
Approximation
$r_{B}=\mu/2$,
$\sigma^{2}_{B}>0.1$
@@ III Ground Truth Calculation
In order to evaluate the performance of the detectors against ideal results, an optimal or “ground truth”(GT) anomaly detector is required. Here, the GT anomaly detector is constructed based on the Bayes Classifier, which can only be used when the class probability distributions are known. Given the probability distribution function for the healthy data is $\mbox{PDF}_{H}$ and faulty is $\mbox{PDF}_{F}$, with data point $x$, the predicted label by the Bayes Classifier is:
@@ IV Experiment
Number of Training Samples
$1000\pm 2$,
$2000\pm 2$,
$5000\pm 2$,
$10000\pm 2$
@@
Number of Simulations
per Sample Size
per Distribution Scenario
100
@@
Scenario
Distribution
Parameters
Ground Truth Metrics
$S_{1}$
$d=2$,
$\sigma^{2}_{A}=0.05$,
$\sigma^{2}_{B}=0.4$,
$\mu=2.8$
$\mbox{FPR}=1.0\%$,
$\mbox{FNR}=7.4\%$,
$\mbox{AUCROC}=99\%$
$S_{2}$
$d=10$,
$\sigma^{2}_{A}=0.02$,
$\sigma^{2}_{B}=0.04$,
$\mu=1.05$
$\mbox{FPR}=1.0\%$,
$\mbox{FNR}=15\%$,
$\mbox{AUCROC}=99\%$
@@ IV-A Test Cases
IForest*
contamination: 0.01
random_state: 0 to 49
n_estimators: 50, 75, 100
max_samples: ’auto’, 0.5, 0.7, 0.8, 0.9
kNN
contamination: 0.01
n_neighbors: 3, 5, 7, 0.001N, 0.01N, 0.02N,
          0.03N, 0.04N, 0.06N, 0.08N,
          0.10N, 0.12N, 0.15N $\dagger$
LOF
contamination: 0.01
n_neighbors: 3, 5, 7, 0.001N, 0.01N, 0.02N,
          0.03N, 0.04N, 0.06N, 0.08N,
          0.10N, 0.12N, 0.15N $\dagger$
OCSVM
contamination: 0.01
kernel: sigmoid, rbf, linear
gamma: auto, scale
nu: 0.3, 0.5, 0.7, 0.9
@@
XGBOD None
random_state: 1, 2, 3, 4, 5
@@
SVM
class_weight: balanced
break_ties: True
kernel: sigmoid, rbf, linear
gamma: auto, scale
c: 0.1, 0.3, 0.5, 0.7, 1.0, 1.2, 1.5, 1.7, 2.0
@@ V-A Test Performance
Total Number of
Training Examples
@@ 80 FS FS FS
Total Number of
Training Examples
@@ 80 $S$ $S$ $S$
Top Five
Detectors
XGBOD (28%) LOF (22%) KNN (13%) CBLOF (11%) FTTransformer (8%) LOF (33%) CBLOF (17%) DeepSAD (16%) KNN (15%) XGBOD (9%) LOF (27%) KNN (20%) SVM (13%) CBLOF (11%) XGBOD (9%) SVM (22%) KNN (18%) LOF (14%) CatB (13%) XGBOD (9%)
Top Five
Detectors
CBLOF (37%) LOF (31%) KNN (19%) IForest (6%) CatB (6%) LOF (45%) KNN (42%) SVM (7%) DeepSAD (3%) XGBOD (1%) SVM (33%) KNN (26%) LOF (24%) CatB (13%) DeepSAD (3%) SVM (57%) CatB (25%) LOF (8%) KNN (6%) XGBOD (3%)
@@
Top Five
Detectors
LOF (25%) KNN (22%) CBLOF (15%) XGBOD (14%) DeepSAD (9%) LOF (36%) KNN (29%) XGBOD (12%) CBLOF (11%) DeepSAD (5%) XGBOD (37%) LOF (26%) KNN (21%) DeepSAD (13%) CBLOF (3%) XGBOD (54%) LOF (19%) KNN (9%) DeepSAD (8%) CBLOF (7%)
Top Five
Detectors
LOF (42%) KNN (35%) CBLOF (22%) XGBOD (0%) LOF (53%) KNN (44%) XGBOD (3%) LOF (48%) KNN (26%) XGBOD (26%) CBLOF (0%) XGBOD (80%) LOF (9%) KNN (7%) CBLOF (3%) SVM (2%)
@@
Table IX: Top five most frequently selected top detectors by validation and test AUCROC across different training dataset sizes (minimum of 199 simulations per sample size). Percent frequency of detector selection is in brackets, rounded to nearest percent. Training anomaly rate is 0.5%.

### v2

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I-A4 Summary
Category Source Name Year Number of Examples Anomaly Rate
@@ II-B Available Datasets
Class Distribution Type Parameters
Healthy Three-mode Gaussian Mixture $\sigma^{2}_{A}>0.01$, $\mu>0$, $\mu_{A}=\sqrt{\mu^{2}/d}$, $\mu_{1}=\vec{0}$, $\mu_{2}=\mu_{A}\vec{1}$, $\mu_{3}=-\mu_{A}\vec{1}$
Faulty Gaussian Mixture Hypersphere Approximation $r_{B}=\mu/2$, $\sigma^{2}_{B}>0.1$
@@ III Ground Truth Calculation
In order to evaluate the performance of the detectors against ideal results, an optimal or “ground truth”(GT) anomaly detector is required. Here, the GT anomaly detector is constructed based on the Bayes Classifier, which can only be used when the class probability distributions are known. Given the probability distribution function for the healthy data is $\mbox{PDF}_{H}$ and faulty is $\mbox{PDF}_{F}$, with data point $x$, the predicted label by the Bayes Classifier is [25]:
@@ IV Experiment
Number of Training Samples $1000\pm 2$, $2000\pm 2$, $5000\pm 2$, $10000\pm 2$
@@
Number of Simulations per Sample Size per Distribution Scenario 100
@@
Scenario Distribution Parameters Ground Truth Metrics
$S_{1}$ $d=2$, $\mu=2.8$ $\sigma^{2}_{A}=0.05$, $\sigma^{2}_{B}=0.4$ $\mbox{FPR}=1.0\%$, $\mbox{FNR}=7.4\%$, $\mbox{AUCROC}=99\%$
$S_{2}$ $d=10$, $\mu=1.05$ $\sigma^{2}_{A}=0.02$, $\sigma^{2}_{B}=0.04$ $\mbox{FPR}=1.0\%$, $\mbox{FNR}=15\%$, $\mbox{AUCROC}=99\%$
@@ IV-A Test Cases
IForest* contamination: 0.01 random_state: 0 to 49 n_estimators: 50, 75, 100 max_samples: ’auto’, 0.5, 0.7, 0.8, 0.9
kNN contamination: 0.01 n_neighbors: 3, 5, 7, 0.001N, 0.01N, 0.02N,           0.03N, 0.04N, 0.06N, 0.08N,           0.10N, 0.12N, 0.15N $\dagger$
LOF contamination: 0.01 n_neighbors: 3, 5, 7, 0.001N, 0.01N, 0.02N,           0.03N, 0.04N, 0.06N, 0.08N,           0.10N, 0.12N, 0.15N $\dagger$
OCSVM contamination: 0.01 kernel: sigmoid, rbf, linear gamma: auto, scale nu: 0.3, 0.5, 0.7, 0.9
@@
XGBOD None random_state: 1, 2, 3, 4, 5
@@
SVM class_weight: balanced break_ties: True kernel: sigmoid, rbf, linear gamma: auto, scale c: 0.1, 0.3, 0.5, 0.7, 1.0, 1.2, 1.5, 1.7, 2.0
@@ V-A Test Performance
Total Number of Training Examples
@@ 80 FS FS FS
Total Number of Training Examples
@@ 80 $S$ $S$ $S$
Top Five Detectors (Validation) XGBOD (28%) LOF (22%) KNN (13%) CBLOF (11%) FTTransformer (8%) LOF (33%) CBLOF (17%) DeepSAD (16%) KNN (15%) XGBOD (9%) LOF (27%) KNN (20%) SVM (13%) CBLOF (11%) XGBOD (9%) SVM (22%) KNN (18%) LOF (14%) CatB (13%) XGBOD (9%)
Top Five Detectors (Test) CBLOF (37%) LOF (31%) KNN (19%) IForest (6%) CatB (6%) LOF (45%) KNN (42%) SVM (7%) DeepSAD (3%) XGBOD (1%) SVM (33%) KNN (26%) LOF (24%) CatB (13%) DeepSAD (3%) SVM (57%) CatB (25%) LOF (8%) KNN (6%) XGBOD (3%)
@@
Top Five Detectors (Validation) LOF (25%) KNN (22%) CBLOF (15%) XGBOD (14%) DeepSAD (9%) LOF (36%) KNN (29%) XGBOD (12%) CBLOF (11%) DeepSAD (5%) XGBOD (37%) LOF (26%) KNN (21%) DeepSAD (13%) CBLOF (3%) XGBOD (54%) LOF (19%) KNN (9%) DeepSAD (8%) CBLOF (7%)
Top Five Detectors (Test) LOF (42%) KNN (35%) CBLOF (22%) XGBOD (0%) LOF (53%) KNN (44%) XGBOD (3%) LOF (48%) KNN (26%) XGBOD (26%) CBLOF (0%) XGBOD (80%) LOF (9%) KNN (7%) CBLOF (3%) SVM (2%)
@@
Table IX: Top five most frequently selected top detectors by validation and test AUCROC across different training dataset sizes (minimum of 199 simulations per sample size, thus 0% may be shown due to rounding). Percent frequency of detector selection is in brackets, rounded to nearest percent. Training anomaly rate is 0.5%.